# Efficient and Universally Accessible Cross-Chain Options without Upfront Holder Collateral

### v2

@@ VII RELATED WORK
Currently, many on-chain options trading platforms are available in the market. They price the options using an Automatic Market Maker (AMM). Lyra [4] stands as the preeminent decentralized options trading platform, commanding around a third of the market’s TVL, and employs an AMM with a Black76 [37] pricing model. However, its operations hinge on external data feeds from oracle, such as spot prices and implied volatility. Hegic [6] decentralizes the writers’ risk and employs a fixed pricing rate based on option expiry date and target prices, which leads to less accurate pricing.
@@
In traditional markets, the price of options is determined by supply and demand. Devising an effective pricing model for options by AMM faces challenges due to the lack of accurate supply and demand modeling. Therefore, an order-book based decentralized exchange shows up. Aevo [7] is a high-performance, order-book based decentralized exchange, which closely resembles the traditional options market. However, its current implementation employs an off-chain orderbook coupled with on-chain settlement, which introduces a higher degree of centralized risk into the protocol. Opyn [5] provides users with the ability to sell European options by minting ERC20 tokens as the option. These tokens can be destroyed to exercise rights or transacted in the market. However, the system faces challenges due to high gas fees on Ethereum and a lack of necessary liquidity for exchanges. All of above on-chain protocols lack universality, most [4, 5, 6, 7] currently only support ETH and BTC options trading.
@@
A HTLC can address the aforementioned issues by enabling two parties to create option contracts across two chains. These contracts lock assets, agreed upon by both parties, on two chains at a predetermined price. HTLCs [38, 24, 39] were originally designed for cross-chain atomic swaps. Subsequently, Han et al. [26] highlighted the optionality and fairness aspects for one party, demonstrating that an atomic cross-chain swap is equivalent to a premium-free American call option. They estimate premiums with Cox-Ross-Rubinstein option pricing model [40]. They addresses the unfairness by incorporating a premium mechanism. In [41], the authors define a sore loser attack in cross-chain swaps and let participants escrow assets along with a negotiated option premium, which acts as compensation. Nadahalli et al. [42] separate the premium protocol from the collateral protocol, employs upfront communication of off-chain unspent transaction outputs as the option premium and collateral. Other researchers [29, 28] introduce cross-chain atomic options, incorporating concepts such as the holder’s late margin deposit and early cancellation of the option. In [30], the authors introduce transferability of options. However, their approach requires long transfer times and does not support concurrent trading involving multiple buyers, which may lead to phantom bid attacks. An adversary can create multiple fake buyers who offer higher prices but fail to complete the transfer. Consequently, the option holder is unable to sell their position.
@@
None of these protocols eliminates the holder’s collateral in cross-chain options. To eliminate the holder’s upfront collateral requirement, cross-chain transaction confirmation can be adopted to verify the collateral deposition on one chain when the option is exercised. This approach can employ cross-chain bridges. Some cross-chain bridges rely on external verification and introduce a trusted third party to facilitate message transmission. This approach is vulnerable to many attacks [12, 9], such as rug pulls [13], code vulnerabilities [14] and private key leakage [15]. Some bridges employ native verification and use light clients on both chains to verify proofs. This method requires complex smart contracts and incurs high verification and storage costs [19, 20, 21, 22]. The diversity and heterogeneity of blockchains significantly increase the time and cost of implementing a light client for each chain. An alternative is using TEE for cross-chain transactions [43, 44]. Those solutions are susceptible to many vulnerabilities, including side-channel attacks, which present significant security risks [16, 17, 18, 45].
@@ VIII CONCLUSION AND DISCUSSION
In this paper, we proposed an approach for efficient and collateral-free cross-chain option. Compared to an HTLC-based option, our approach incorporates a signature scheme DAPS to facilitate efficient option transfer, reducing the transfer time to half of that of previous work. Akin to traditional options, our approach does not require the option holder to put a collateral upfront when an option is established, thus make the most of the advantages of options for leveraging and hedging. With our proposed method, the option markets can be more efficient.
One key feature of our proposed protocol is its resilience against phantom bidder attacks. The protocol supports concurrent bidding and can finalize the option transfer in a constant time, not influenced by potential malicious option buyers.
In practice, we can further utilize the liquidity of the writer’s collateral. For example, Bob, the option writer, can still use the funds in the contract for trading, anyone can flash loan Bob’s collateral completed in a single transaction.
The options we explored in this paper are actually covered options. In option markets, there are other types of options, for example, naked options where sellers must maintain their margin accounts at a threshold according to market fluctuations, otherwise face liquidation. In the future, we are interested in designing naked options with liquidation features and consider how to ensure the rapid transfer of cross-chain (naked) options.

### v4

@@ VII RELATED WORK
Currently, many on-chain options trading platforms are available in the market. They price the options using an Automatic Market Maker (AMM). Lyra [4] stands as the preeminent decentralized options trading platform, commanding around a third of the market’s TVL, and employs an AMM with a Black76 [37] pricing model. However, its operations hinge on external data feeds from oracles, such as spot prices and implied volatility. Hegic [6] decentralizes the writers’ risk and employs a fixed pricing rate based on option expiry date and target prices, which leads to less accurate pricing.
@@
In traditional markets, the price of options is determined by supply and demand. Devising an effective pricing model for options by AMM faces challenges due to the lack of accurate supply and demand modeling. Therefore, an order-book-based decentralized exchange shows up. Aevo [7] is a high-performance, order-book-based decentralized exchange, which closely resembles the traditional options market. However, its current implementation employs an off-chain orderbook coupled with on-chain settlement, which introduces a higher degree of centralized risk into the protocol. Opyn [5] provides users with the ability to sell European options by minting ERC20 tokens as the option. These tokens can be destroyed to exercise rights or transacted in the market. However, the system faces challenges due to high gas fees on Ethereum and a lack of necessary liquidity for exchanges. All of the above on-chain protocols lack universality, most [4, 5, 6, 7] currently only support ETH and BTC options trading.
@@
An HTLC can address the aforementioned issues by enabling two parties to create option contracts across two chains. These contracts lock assets, agreed upon by both parties, on two chains at a predetermined price. HTLCs [38, 24, 39] were originally designed for cross-chain atomic swaps. Subsequently, Han et al. [26] highlighted the optionality and fairness aspects for one party, demonstrating that an atomic cross-chain swap is equivalent to a premium-free American call option. They estimate premiums with the Cox-Ross-Rubinstein option pricing model [40]. They address the unfairness by incorporating a premium mechanism. In [41], the authors define a sore loser attack in cross-chain swaps and let participants escrow assets along with a negotiated option premium, which acts as compensation. Nadahalli et al. [42] separate the premium protocol from the collateral protocol, employing upfront communication of off-chain unspent transaction outputs as the option premium and collateral. Other researchers [29, 28] introduce cross-chain atomic options, incorporating concepts such as the holder’s late margin deposit and early cancellation of the option.
In TCO [30], the authors introduce the transferability of options. However, their approach requires long transfer times and does not support concurrent trading involving multiple buyers, which may lead to phantom bid attacks. An adversary can create multiple fake buyers who offer higher prices but fail to complete the transfer. Consequently, the option holder is unable to sell their position.
@@
None of these protocols eliminates the holder’s collateral in cross-chain options. To eliminate the holder’s upfront collateral requirement, cross-chain transaction confirmation can be adopted to verify the collateral deposition on one chain when the option is exercised. This approach can employ cross-chain bridges. Some cross-chain bridges rely on external verification and introduce a trusted third party to facilitate message transmission. This approach is vulnerable to many attacks [12, 9], such as rug pulls [13], code vulnerabilities [14], and private key leakage [15]. Some bridges employ native verification and use light clients on both chains to verify proofs. This method requires complex smart contracts and incurs high verification and storage costs [19, 20, 21, 22]. The diversity and heterogeneity of blockchains significantly increase the time and cost of implementing a light client for each chain. An alternative is using TEE for cross-chain transactions [43, 44]. Those solutions are susceptible to many vulnerabilities, including side-channel attacks, which present significant security risks [16, 17, 18, 45].
@@ VIII CONCLUSION AND DISCUSSION
In this paper, we propose an approach for efficient and collateral-free cross-chain options. Compared to an HTLC-based option, our approach incorporates a signature scheme, DAPS, to facilitate efficient option transfer, reducing the transfer time to half of the previous work. Akin to traditional options, our approach does not require the option holder to put a collateral upfront when an option is established, thus making the most of the advantages of options for leveraging and hedging. With our proposed method, the option markets can be more efficient.
One key feature of our proposed protocol is its resilience against phantom bidder attacks. The protocol supports concurrent bidding and can finalize the option transfer in constant time, not influenced by potential malicious option buyers.
In practice, we can further utilize the liquidity of the writer’s collateral. For example, Bob, the option writer, can still use the funds in the contract for trading, anyone can flash loan Bob’s collateral, completed in a single transaction.
The options we explored in this paper are covered options. In option markets, there are other types of options, for example, naked options, where sellers must maintain their margin accounts at a threshold according to market fluctuations, otherwise face liquidation. In the future, we plan to design naked options with liquidation features and consider how to ensure the rapid transfer of cross-chain options.